# RETR: Multi-View Radar Detection Transformer for Indoor Perception

### v1

@@ Abstract
Indoor radar perception has seen rising interest due to affordable costs driven by emerging automotive imaging radar developments and the benefits of reduced privacy concerns and reliability under hazardous conditions (e.g., fire and smoke). However, existing radar perception pipelines fail to account for distinctive characteristics of the multi-view radar setting. In this paper, we propose Radar dEtection TRansformer (RETR), an extension of the popular DETR architecture, tailored for multi-view radar perception. RETR inherits the advantages of DETR, eliminating the need for hand-crafted components for object detection and segmentation in the image plane. More importantly, RETR incorporates carefully designed modifications such as 1) depth-prioritized feature similarity via a tunable positional encoding (TPE); 2) a tri-plane loss from both radar and camera coordinates; and 3) a learnable radar-to-camera transformation via reparameterization, to account for the unique multi-view radar setting. Evaluated on two indoor radar perception datasets, our approach outperforms existing state-of-the-art methods by a margin of $15.38+$ AP for object detection and $11.77+$ IoU for instance segmentation, respectively.
@@ 1 Introduction
    Extending DETR for Multi-View Radar Perception: 1) Encoder: we associate features from both radar views by applying self-attention over the pooled multi-view radar tokens, eliminating the need for a cumbersome association scheme. We introduce a top-$K$ feature selection to allow only $K$ features from each view to keep the complexity low. 2) Decoder: the DETR decoder provides a natural way to associate the same object query to corresponding features from the two radar views via cross-attention. As such, the object query is able to learn 3D spatial embedding of objects in the radar coordinate (see Fig. 1).
@@ 4.1 RETR Architecture
Tokenization:
A transformer-based encoder expects a sequence of tokens as input. This is done by mapping the feature maps into a sequence of $P$ multi-view radar tokens $\bm{H}=\left\{{\bm{H}_{\mathtt{hor}},\bm{H}_{\mathtt{ver}}}\right\}\in\mathbb{R}^{C\times P}$: $\bm{Z}_{\mathtt{hor}}\rightarrow\bm{H}_{\mathtt{hor}}\in\mathbb{R}^{C\times P_{\mathtt{hor}}}$ and $\bm{Z}_{\mathtt{ver}}\rightarrow\bm{H}_{\mathtt{ver}}\in\mathbb{R}^{C\times P_{\mathtt{ver}}}$, where $P=P_{\mathtt{hor}}+P_{\mathtt{ver}}$. We defer the tokenization discussion to Section 4.2.
@@
The transformer encoder provides a simple yet effective method for associating radar features from both horizontal and vertical views by applying self-attention over the pool of $P$ multi-view radar tokens $\bm{H}=\left\{{\bm{H}_{\mathtt{hor}},\bm{H}_{\mathtt{ver}}}\right\}\in\mathbb{R}^{C\times P}$, eliminating the need for cumbersome association schemes. Specifically, the $l$-th ($l=0,\cdots,L_{\mathtt{self}}-1$) encoder layer updates the multi-view radar tokens through multi-head self-attention $\mathtt{Att}_{\mathtt{self}}$:
@@
Additionally, since the multi-view radar tokens lack positional information and the self-attention is permutation-invariant, we supplement $\bm{H}^{l}$ with positional embedding added (or attached) to the input of each encoder layer. Refer to Section 4.3 for a tunable positional encoding.
@@
where $\mathtt{Que}$, $\mathtt{Key}$ and $\mathtt{Val}$ are the projections with different parameterization from those in the self-attention layer (Eq. 3). Then, the decoder layer further updates the object queries $\bar{\bm{Q}}^{l}$ of Eq. 4 via multi-head cross-attention with the multi-view radar tokens $\bm{H}^{L_{\mathtt{self}}}$ from the encoder output:
@@
4.2 Top-$K$ Feature Selection as Tokenization
In DETR, the tokenization simply collapses the spatial dimensions of the feature map into a single dimension, resulting in $P_{\mathtt{hor}}=WD/s^{2}$ and $P_{\mathtt{ver}}=HD/s^{2}$ tokens for the horizontal and vertical radar feature maps, respectively. As a result, we have $P=(W+H)D/s^{2}$ multi-view radar tokens. It is known that the complexity of transformers grows quadratically with respect to the token length $P$. Here, we introduce a customized Top-$K$ feature selection as tokenization, maintaining a low complexity for the RETR encoder and decoder: $\bm{H}_{\mathtt{hor}}=\mathtt{Selector}\left({\bm{Z}_{\mathtt{hor}}}\right)\in\mathbb{R}^{C\times K}$ and $\bm{H}_{\mathtt{ver}}=\mathtt{Selector}\left({\bm{Z}_{\mathtt{ver}}}\right)\in\mathbb{R}^{C\times K}$, where $K\ll\min\{WD/s^{2},HD/s^{2}\}$. In this case, we shrink the multi-view radar tokens from $P=(W+H)D/s^{2}$ to $P=2K$. For each radar frame, we consistently select the Top-$K$ strongest features, which may originate from varying spatial locations depending on the specific radar frame. Consequently, the gradient propagates back through the selected $K$ features to the backbone weights, irrelevant to their spatial locations.
@@ Appendix B Segmentation
RETR 3D $\mathtt{H}$, $\mathtt{V}$ $\mathtt{H}+\mathtt{V}+\mathtt{I}$ 77.07
@@ Appendix G Additional Ablation Study
RETR 3D $\mathtt{H}$, $\mathtt{V}$ $\mathtt{H}+\mathtt{V}+\mathtt{I}$ 12.45 41.30 4.96 19.96 21.58
@@
$\mathtt{H}+\mathtt{V}+\mathtt{I}$ 46.75 83.80 46.06 42.19 57.39 77.07

### v3

@@ Abstract
Indoor radar perception has seen rising interest due to affordable costs driven by emerging automotive imaging radar developments and the benefits of reduced privacy concerns and reliability under hazardous conditions (e.g., fire and smoke). However, existing radar perception pipelines fail to account for distinctive characteristics of the multi-view radar setting. In this paper, we propose Radar dEtection TRansformer (RETR), an extension of the popular DETR architecture, tailored for multi-view radar perception. RETR inherits the advantages of DETR, eliminating the need for hand-crafted components for object detection and segmentation in the image plane. More importantly, RETR incorporates carefully designed modifications such as 1) depth-prioritized feature similarity via a tunable positional encoding (TPE); 2) a tri-plane loss from both radar and camera coordinates; and 3) a learnable radar-to-camera transformation via reparameterization, to account for the unique multi-view radar setting. Evaluated on two indoor radar perception datasets, our approach outperforms existing state-of-the-art methods by a margin of $15.38+$ AP for object detection and $11.91+$ IoU for instance segmentation, respectively. Our implementation is available at https://github.com/merlresearch/radar-detection-transformer.
@@ 1 Introduction
    Extending DETR for Multi-View Radar Perception: 1) Encoder: we associate features from both radar views by applying self-attention over the pooled multi-view radar features, eliminating the need for a cumbersome association scheme. We introduce a top-$K$ feature selection to allow only $K$ features from each view to keep the complexity low. 2) Decoder: the DETR decoder provides a natural way to associate the same object query to corresponding features from the two radar views via cross-attention. As such, the object query is able to learn 3D spatial embedding of objects in the radar coordinate (see Fig. 1).
@@ 4.1 RETR Architecture
Feature Selection:
A transformer-based encoder expects a sequence of features as input. This is done by mapping the feature maps into a sequence of $P$ multi-view radar features $\bm{H}=\left\{{\bm{H}_{\mathtt{hor}},\bm{H}_{\mathtt{ver}}}\right\}\in\mathbb{R}^{C\times P}$: $\bm{Z}_{\mathtt{hor}}\rightarrow\bm{H}_{\mathtt{hor}}\in\mathbb{R}^{C\times P_{\mathtt{hor}}}$ and $\bm{Z}_{\mathtt{ver}}\rightarrow\bm{H}_{\mathtt{ver}}\in\mathbb{R}^{C\times P_{\mathtt{ver}}}$, where $P=P_{\mathtt{hor}}+P_{\mathtt{ver}}$. We defer the discussion of top-$K$ feature selection to Section 4.2.
@@
The transformer encoder provides a simple yet effective method for associating radar features from both horizontal and vertical views by applying self-attention over the pool of $P$ multi-view radar features $\bm{H}=\left\{{\bm{H}_{\mathtt{hor}},\bm{H}_{\mathtt{ver}}}\right\}\in\mathbb{R}^{C\times P}$, eliminating the need for cumbersome association schemes. Specifically, the $l$-th ($l=0,\cdots,L_{\mathtt{self}}-1$) encoder layer updates the multi-view radar features through multi-head self-attention $\mathtt{Att}_{\mathtt{self}}$:
@@
Additionally, since the multi-view radar features lack positional information and the self-attention is permutation-invariant, we supplement $\bm{H}^{l}$ with positional embedding added (or attached) to the input of each encoder layer. Refer to Section 4.3 for a tunable positional encoding.
@@
where $\mathtt{Que}$, $\mathtt{Key}$ and $\mathtt{Val}$ are the projections with different parameterization from those in the self-attention layer (Eq. 3). Then, the decoder layer further updates the object queries $\bar{\bm{Q}}^{l}$ of Eq. 4 via multi-head cross-attention with the multi-view radar features $\bm{H}^{L_{\mathtt{self}}}$ from the encoder output:
@@
4.2 Top-$K$ Feature Selection
In DETR, the sequentialization simply collapses the spatial dimensions of the feature map into a single dimension, resulting in $P_{\mathtt{hor}}=WD/s^{2}$ and $P_{\mathtt{ver}}=HD/s^{2}$ features for the horizontal and vertical radar feature maps, respectively. As a result, we have $P=(W+H)D/s^{2}$ multi-view radar features. It is known that the complexity of transformers grows quadratically with respect to the feature length $P$. Here, we introduce a customized Top-$K$ feature selection, maintaining a low complexity for the RETR encoder and decoder: $\bm{H}_{\mathtt{hor}}=\mathtt{Selector}\left({\bm{Z}_{\mathtt{hor}}}\right)\in\mathbb{R}^{C\times K}$ and $\bm{H}_{\mathtt{ver}}=\mathtt{Selector}\left({\bm{Z}_{\mathtt{ver}}}\right)\in\mathbb{R}^{C\times K}$, where $K\ll\min\{WD/s^{2},HD/s^{2}\}$. In this case, we shrink the multi-view radar tokens from $P=(W+H)D/s^{2}$ to $P=2K$. For each radar frame, we consistently select the Top-$K$ strongest features, which may originate from varying spatial locations depending on the specific radar frame. Consequently, the gradient propagates back through the selected $K$ features to the backbone weights, irrelevant to their spatial locations.
@@ Appendix B Segmentation
RETR 3D $\mathtt{H}$, $\mathtt{V}$ $\mathtt{H}+\mathtt{V}+\mathtt{I}$ 77.21
@@ Appendix G Additional Ablation Study
RETR 3D $\mathtt{H}$, $\mathtt{V}$ $\mathtt{H}+\mathtt{V}+\mathtt{I}$ 12.19 40.67 4.95 19.70 21.34
@@
$\mathtt{H}+\mathtt{V}+\mathtt{I}$ 46.75 83.80 46.06 42.19 57.39 77.21